Think gravity battery. 

  1. How much suspended weight and how far down would it need to be lowered in order to produce 100kwh? 
  2. How many Kw would it take to lift the weight back up? 
  3. What style of motor/generator can do this? It's being used both as the generator and the motor.

A direct drive AC high torque would be good to keep the RPM's and maintenance low. Any experts on motors here? 


1
Expert's answer
2021-04-12T01:54:12-0400

To achieve 100 kW, you would need to lower a certain weight of mass m for certain height h during certain time t:


"P=\frac{E}{t}=\frac{mgh}{t}."

This equation ties these three values together. If you want to use a motor as a generator, it is hard to tell which model is better because motors are designed to work as motors, not generators, although they can produce energy.

You will need more energy to lift this weight back up. Only a proper experiment would shed light on it.

Imagine you want to lift 900 kg for 50 m with a 100-kW motor. This will take a time of


"t=\frac{mgh}{P}=4.41 \text{ s}."

DC motors can work as generators. Synchronous AC motors can work as generators.